# Multifidelity Simulation-based Inference for Computationally Expensive Simulators

### v3

@@ Abstract
Across many domains of science, stochastic models are an essential tool to understand the mechanisms underlying empirically observed data. Models can be of different levels of detail and accuracy, with models of high-fidelity (i.e., high accuracy) to the phenomena under study being often preferable. However, inferring parameters of high-fidelity models via simulation-based inference is challenging, especially when the simulator is computationally expensive. We introduce MF-(TS)NPE, a multifidelity approach to neural posterior estimation that uses transfer learning to leverage inexpensive low-fidelity simulations to efficiently infer parameters of high-fidelity simulators. MF-(TS)NPE applies the multifidelity scheme to both amortized and non-amortized neural posterior estimation. We further improve simulation efficiency by introducing A-MF-TSNPE, a sequential variant that uses an acquisition function targeting the predictive uncertainty of the density estimator to adaptively select high-fidelity parameters. On established benchmark and neuroscience tasks, our approaches require up to two orders of magnitude fewer high-fidelity simulations than current methods, while showing comparable performance. Overall, our approaches open new opportunities to perform efficient Bayesian inference on computationally expensive simulators.
@@ 1 Introduction
Stochastic models are used across science and engineering to capture complex properties of real systems through simulations (Barbers et al., 2024; Nelson & Pei, 2021; Fadikar et al., 2018; Pillow & Scott, 2012). These simulators encode domain-specific knowledge and provide a means to generate high-fidelity synthetic data, enabling accurate forward modeling of experimental outcomes. However, inferring model parameters from observed data can be challenging, especially when simulators are stochastic, the likelihoods of the simulators are inaccessible, or when simulations are computationally expensive.
Bayesian inference offers a principled framework to address the challenge of estimating parameters from stochastic observations by quantifying uncertainty in parameter estimates while incorporating prior knowledge. However, classical Bayesian methods are not applicable to cases where the likelihood of the simulator is not accessible in closed form (Tavaré et al., 1997). Simulation-based inference (SBI) has emerged as a paradigm to tackle this by leveraging forward simulations to infer the posterior distribution (Cranmer et al., 2020).
@@
The challenge of extending sampling-based SBI methods like Approximate Bayesian Computation (ABC) (Tavaré et al., 1997; Pritchard et al., 1999) to problems with large numbers of parameters has driven significant advancements (Cranmer et al., 2020). Modern SBI approaches aim for the estimation of the posterior through amortized neural posterior estimation (NPE) (Greenberg et al., 2019; Lueckmann et al., 2017; Papamakarios & Murray, 2016), neural likelihood estimation (Papamakarios et al., 2019), neural ratio estimation (Hermans et al., 2020; Thomas et al., 2022) and respective sequential variants. These approaches have leveraged recent progress in neural density estimation to improve the scalability and accuracy of SBI, allowing parameter inference in problems with higher dimensionality than was previously achievable (Ramesh et al., 2021; Gloeckler et al., 2024). Despite these advancements, SBI methods face computational challenges for scenarios involving expensive simulations or high-dimensional parameter spaces, as state-of-the-art methods often require extensive simulation budgets to achieve reliable posterior estimates (Lueckmann et al., 2021).
@@
Multifidelity modeling offers a solution to this problem by balancing precision and efficiency. It combines accurate but costly high-fidelity models (Hoppe et al., 2021; Behrens & Dias, 2015) with faster, less accurate low-fidelity models. Here, low-fidelity models could be simplifications made possible through domain knowledge about the high-fidelity models, low-dimensional projection of the high-fidelity model, or surrogate modeling (Peherstorfer et al., 2018). For example, Reynolds-averaged Navier-Stokes (RANS) models simplify turbulent flow simulations in aerodynamics (Han et al., 2013), while climate models often reduce complexity by focusing on specific atmospheric effects (Held, 2005; Majda & Gershgorin, 2010). Similarly, mean-field approximations are used to capture certain features of spiking neural network dynamics (Vogels et al., 2011; Dayan & Abbott, 2001). By exploiting this range of fidelities, recent methods in inference, such as multifidelity Monte Carlo approaches (Peherstorfer et al., 2016; Nobile & Tesei, 2015; Giles, 2008; Zeng et al., 2023) have shown clear improvements in computational efficiency through the incorporation of coarse approximations as low-fidelity models. In the context of SBI, we hypothesized that by leveraging the complementarity of high- and low-fidelity simulators, it would be possible to reduce the computational cost of inference while retaining inference accuracy.
@@
In this work, we present MF-(TS)NPE, a multifidelity approach that improves the efficiency of amortized neural posterior estimation for expensive simulators. MF-(TS)NPE reduces the computational burden of posterior estimation by pre-training a neural density estimator on low-fidelity simulations and refining the inference with a smaller set of high-fidelity simulations. Additionally, we present A-MF-TSNPE, an extension of MF-TSNPE with active learning, facilitating targeted parameter space exploration to effectively enhance high-fidelity posterior estimates given single observations. We focus on multifidelity cases where both models are simulators and where the low-fidelity model is a simplified version of the high-fidelity model, designed based on domain expertise. We demonstrate that for four benchmark tasks and two computationally expensive neuroscience simulators, our multifidelity approach can identify the posterior distributions more efficiently than NPE and TSNPE, often reducing the number of required high-fidelity simulations by orders of magnitude.
@@ 2 Background
Multifidelity has been widely explored in the context of likelihood-based inference (Peherstorfer et al., 2018), from maximum likelihood estimation approaches (Maurais et al., 2023) to Bayesian inference methods (Vo et al., 2019; Catanach et al., 2020). For cases where the likelihood is not explicitly available, several sampling-based multifidelity methods have been proposed within the framework of ABC (Prescott & Baker, 2020; Warne et al., 2022; Prescott et al., 2024; Prescott & Baker, 2021). However, these methods inherit limitations of ABC approaches, particularly in high-dimensional parameter spaces, where neural density estimators offer more scalable alternatives to complex real-world problems (Lueckmann et al., 2021). Concurrently with our work, Thiele et al. (2025) developed a multifidelity SBI approach based on response distillation, Hikida et al. (2025) adapted multilevel Monte Carlo techniques to SBI, and Saoulis et al. (2025) applied transfer learning to accelerate inference on a cosmological task.
Beyond SBI, multifidelity has been explored in Bayesian optimization, where Gaussian process models integrate data of different fidelities to infer expensive functions (e.g., Song et al., 2019). These approaches focus on learning surrogate likelihood functions rather than posteriors over simulator parameters, but they highlight the broad applicability of the multifidelity concept.
@@
To facilitate learning in a target domain, transfer learning borrows knowledge from a source domain (Panigrahi et al., 2021). This is often done when the target dataset is smaller than the source dataset (Larsen-Freeman, 2013) and has successfully been applied to a range of machine learning tasks, e.g., in computer vision (Jiang & Learned-Miller, 2017; Hussain et al., 2019). In the context of numerical simulators, transfer learning approaches have been used to lower the simulation budget, for instance, in CO2 forecasting (Falola et al., 2023), surrogate modeling (Wang et al., 2024) and model inversion with physics-informed neural networks (Haghighat et al., 2021). To the best of our knowledge, the potential of transfer learning for computationally efficient simulation-based inference has not been fully realized yet.
@@
Recent works have similarly sought to reduce the computational burden of SBI for expensive simulators with active learning frameworks to select informative simulation parameters. Lueckmann et al. (2019) focuses on likelihood estimation, while Griesemer et al. (2024) directly estimates the posterior distribution. Dyer et al. (2022) targets inference under extremely low simulation budgets, employing signature-based feature extraction paired with amortized ratio estimation. (Gutmann & Corander, 2016) proposed an ABC method that uses Bayesian optimisation to efficiently model the discrepancy between the simulator output and the empirical data. More recently, Gloeckler et al. (2025) proposes a compositional framework for time series models that exploits their Markovian structure. Schmitt et al. (2024a) and Schmitt et al. (2024b) introduce self-consistency and consistency models to regularize amortized inference and scale SBI. Delaunoy et al. (2024) show that Bayesian neural networks can be used to obtain calibrated posterior approximations under small simulation budgets. These methods share with MF-NPE the goal of maximizing simulation efficiency through feature extraction and active learning, although they remain within a single-fidelity context.
A related approach involves learning surrogate models to approximate high-fidelity simulators (Wiqvist et al., 2021; Dyer et al., 2023). Both methods train neural surrogate simulators alongside posterior or likelihood estimators, avoiding handcrafted low-fidelity models. In contrast, MF-(TS)NPE assumes the availability of a low-fidelity simulator derived from domain expertise, leveraging transfer learning and active learning strategies to selectively refine posterior estimates.
@@ 3 Methods
MF-(TS)NPE is a multifidelity approach to neural posterior estimation for computationally expensive simulators. We present our approach in Sec. 3.1. In Sec. 3.1.4, we discuss the evaluation metrics used to compare our method against NPE (Greenberg et al., 2019), TSNPE (Deistler et al., 2022), and MF-ABC (Prescott & Baker, 2020). MF-(TS)NPE is summarized in Fig. 1, Algorithms 1 and 3.
@@ 3.1 Multifidelity NPE
Figure 1: Multifidelity Neural Posterior Estimation proceeds by dense sampling from the prior distribution, running the low-fidelity simulator (e.g., a two-compartment neuron model (Hodgkin & Huxley, 1952)), and training a neural density estimator with a negative log-likelihood loss. MF-NPE then retrains the pre-trained network on sparse samples from the same prior distribution and respective high-fidelity simulations (e.g., a multicompartmental neuron model (Rall, 1995)). Given empirical observations $\boldsymbol{x_{o}}$, MF-NPE estimates the posterior distribution given the high-fidelity model. In the sequential case, the parameters for high-fidelity simulations are drawn from iterative refinements of the prior distribution within the support of the current posterior estimate, at some observation $\boldsymbol{x_{o}}$.
@@
As with NPE (Papamakarios & Murray, 2016; Greenberg et al., 2019), to estimate the posterior density over model parameters $\boldsymbol{\theta}$ for which the likelihood function is unavailable, we consider a sufficiently expressive neural density estimator $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})$, and train it to minimize the negative log-likelihood loss:
@@
where $\boldsymbol{\theta}$ is sampled from the prior distribution, $\boldsymbol{x}$ denotes the respective simulations (i.e., samples from $p(\boldsymbol{x}|\boldsymbol{\theta})$), and $\boldsymbol{\phi}$ are the network parameters. By minimizing $\mathcal{L}(\cdot)$, the neural density estimator approximates the conditional distribution $p(\boldsymbol{\theta}|\boldsymbol{x})$ directly (Papamakarios & Murray, 2016) (proof of convergence in Appendix C). Given an empirical observation $\boldsymbol{x_{o}}$, we can then estimate the posterior over parameters $p(\boldsymbol{\theta}|\boldsymbol{x_{o}})$. To ensure $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x_{o}})$ closely approximates the true posterior $p(\boldsymbol{\theta}|\boldsymbol{x_{o}})$, the density estimator must be sufficiently expressive. We use neural spline flows (NSFs) (Durkan et al., 2019), expressive normalizing flows that have been shown empirically to be competitive for SBI (Lueckmann et al., 2021). To avoid overfitting when training NSFs, we use the same validation-based early stopping criterion $S$ as in the SBI package (Boelts et al., 2024) (details in Appendix.D.1).
@@ 3.1.1 Transfer learning
MF-NPE leverages representations learned from low-fidelity simulations to reduce the number of high-fidelity simulations required to approximate a high-fidelity posterior. To that end, MF-NPE adopts a fine-tuning strategy of transfer learning: Let $\psi$ be the parameters of the low-fidelity neural density estimator $q_{\psi}(\boldsymbol{\theta}|\boldsymbol{x}_{\mathrm{L}})$ and let $\phi$ be the parameters of the high-fidelity density estimator $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})$. MF-NPE minimizes the loss $\mathcal{L(\phi)}=\mathbb{E}_{\theta\sim p(\boldsymbol{\theta})}\mathbb{E}_{x\sim p(\boldsymbol{x}|\boldsymbol{\theta})}\left[-\log q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})\right]$ on the high-fidelity task, where the parameters $\phi$ are initialized on the pretrained low-fidelity network parameters $\psi$. We argue that by pre-training on low-fidelity simulations, the density estimator learns useful features up front (i.e., the feature spaces of the low- and high-fidelity density estimators overlap), so fewer high-fidelity simulations suffice to refine the posterior estimates. Indeed, Tahir et al. (2024) shows that once networks learn suitable features for a given predictive task, they drastically reduce the sample complexity for related tasks. Other strategies to pretraining are discussed in Appendix H.3.
MF-NPE can naturally accommodate more than two fidelity levels (Appendix L), does not require more hyperparameter tuning than NPE (Appendix D.1), and is applicable in situations where the low-fidelity model has fewer parameters than the high-fidelity model. In this setting, the parameters that are exclusive to the high-fidelity model are treated as dummy variables in the pre-trained density estimator. The pre-conditioning with these variables leads to the pre-trained neural density estimator to effectively estimate the prior distribution over the respective parameters (OU3 and OU4 in Appendix J.1). As shown below, our method is compatible with both embedding networks and hand-crafted summary statistics of the observations. Overall, MF-NPE shows competitive performance compared to NPE, even when the dimensionality of the low- and high-fidelity models is substantially different (e.g., Sec. 4.3, where the low- and high-fidelity models have 12 and 24 parameters, respectively).
@@
Input: NN pairs of $(\boldsymbol{\theta},\boldsymbol{x}_{\text{L}})$; $M$ pairs of ($\boldsymbol{\theta},\boldsymbol{x})$; conditional density estimators $q_{\psi}(\boldsymbol{\theta}|\boldsymbol{x}_{\text{L}})$ and $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})$ with respectively learnable parameters $\psi$ and $\phi$; early stopping criterion $S$.
$\mathcal{L}(\psi)=\frac{1}{N}\sum_{i=1}^{N}-\log q_{\psi}\left(\boldsymbol{\theta}_{i}|\boldsymbol{x}^{\text{L}}_{i}\right)$ .
for epoch in epochs do
  train $q_{\psi}$ to minimize $\mathcal{L(\psi)}$ until $S$ is reached.
end for
 Initialize $q_{\phi}$ with weights and biases of trained $q_{\psi}$.
$\mathcal{L}(\phi)=\frac{1}{M}\sum_{i=1}^{M}-\log q_{\phi}\left(\boldsymbol{\theta}_{i}|\boldsymbol{x}_{i}\right)$.
for epoch in epochs do
  train $q_{\phi}$ to minimize $\mathcal{L(\phi)}$ until $S$ is reached.
end for

### v4

@@ Abstract
Across many domains of science, stochastic models are an essential tool to understand the mechanisms underlying empirically observed data. Models can be of different levels of detail and accuracy, with models of high-fidelity (i.e., high accuracy) to the phenomena under study being often preferable. However, inferring parameters of high-fidelity models via simulation-based inference is challenging, especially when the simulator is computationally expensive. We introduce a multifidelity approach to neural posterior estimation that uses transfer learning to leverage inexpensive low-fidelity simulations to efficiently infer parameters of high-fidelity simulators. Our method applies the multifidelity scheme to both amortized and non-amortized neural posterior estimation. We further improve simulation efficiency by introducing a sequential variant that uses an acquisition function targeting the predictive uncertainty of the density estimator to adaptively select high-fidelity parameters. On established benchmark and neuroscience tasks, our approaches require up to two orders of magnitude fewer high-fidelity simulations than current methods, while showing comparable performance. Overall, our approaches open new opportunities to perform efficient Bayesian inference on computationally expensive simulators.
@@ 1 Introduction
Stochastic models are used across science and engineering to capture complex properties of real systems through simulations (Barbers et al., 2024; Nelson and Pei, 2021; Pillow and Scott, 2012; Marlier et al., 2021). These simulators encode domain-specific knowledge and provide a means to generate high-fidelity synthetic data, enabling accurate forward modeling of experimental outcomes. However, inferring model parameters from observed data can be challenging, especially when simulators are stochastic, the likelihoods of the simulators are inaccessible, or when simulations are computationally expensive.
@@
Simulation-based inference (SBI) addresses these challenges by leveraging forward simulations to infer the posterior distribution, enabling quantification of uncertainty even when the likelihood is intractable (Cranmer et al., 2020). The challenge of extending sampling-based SBI methods like Approximate Bayesian Computation (ABC) (Tavaré et al., 1997; Pritchard et al., 1999) to problems with large numbers of parameters has driven significant advancements in neural-based approaches that estimate the likelihood (Papamakarios et al., 2019), the likelihood-to-evidence ratio (Hermans et al., 2020), or directly the posterior (Greenberg et al., 2019; Lueckmann et al., 2017; Papamakarios and Murray, 2016). In particular, amortized Neural Posterior Estimation (NPE) trains a neural density estimator to directly approximate the posterior, bypassing the need to estimate the model evidence (Papamakarios and Murray, 2016). To improve inference for a fixed observation and allow stable training, truncated sequential variants have been introduced for neural posterior estimation (TSNPE) (Deistler et al., 2022), and neural ratio estimation (Miller et al., 2021). These approaches have leveraged recent progress in neural density estimation to improve the scalability and accuracy of SBI, allowing parameter inference in problems with higher dimensionality than was previously achievable (Ramesh et al., 2025; Gloeckler et al., 2024). Despite these advancements, SBI methods face computational challenges for scenarios involving expensive simulations or high-dimensional parameter spaces, as state-of-the-art methods often require extensive simulation budgets to achieve reliable posterior estimates (Lueckmann et al., 2021).
@@
Multifidelity modeling offers a solution to this problem by balancing precision and efficiency. It combines accurate but costly high-fidelity models (Hoppe et al., 2021; Behrens and Dias, 2015) with faster, less accurate low-fidelity models. Here, low-fidelity models could be simplifications made possible through domain knowledge about the high-fidelity models, low-dimensional projection of the high-fidelity model, or surrogate modeling (Peherstorfer et al., 2018). For example, Reynolds-averaged Navier-Stokes (RANS) models simplify turbulent flow simulations in aerodynamics (Han et al., 2013), while climate models often reduce complexity by focusing on specific atmospheric effects (Held, 2005; Majda and Gershgorin, 2010). Similarly, mean-field approximations are used to capture certain features of spiking neural network dynamics (Vogels et al., 2011; Dayan and Abbott, 2001). Multifidelity methods have proven effective across domains—enhancing optimization through multifidelity Bayesian optimization (Song et al., 2019; Kandasamy et al., 2017), and improving the efficiency of inference through multifidelity Monte Carlo approaches (Peherstorfer et al., 2016; Nobile and Tesei, 2015; Giles, 2008; Zeng et al., 2023). In the context of SBI, we hypothesized that by leveraging the complementarity of high- and low-fidelity simulators, it would be possible to reduce the computational cost of inference while retaining inference accuracy.
@@
In this work, we present MF-(TS)NPE, a multifidelity approach that improves the efficiency of amortized and non-amortized neural posterior estimation for expensive simulators. MF-(TS)NPE reduces the computational burden of posterior estimation by pre-training a neural density estimator on low-fidelity simulations and refining the inference with a smaller set of high-fidelity simulations. Additionally, we present MF-TSNPE-AF, an extension of MF-TSNPE with active learning, facilitating targeted parameter space exploration to effectively enhance high-fidelity posterior estimates given single observations. We focus on multifidelity cases where both models are simulators and where the low-fidelity model is a simplified version of the high-fidelity model, designed based on domain expertise. We demonstrate that for four benchmark tasks and two computationally expensive neuroscience simulators, our multifidelity approach can identify the posterior distributions more efficiently than NPE and TSNPE, often reducing the number of required high-fidelity simulations by orders of magnitude.
@@ 2 Background
Multifidelity has been widely explored in the context of likelihood-based inference (Peherstorfer et al., 2018), from maximum likelihood estimation approaches (Maurais et al., 2023) to Bayesian inference methods (Vo et al., 2019; Catanach et al., 2020). For cases where the likelihood is not explicitly available, several sampling-based multifidelity methods have been proposed within the framework of ABC (Prescott and Baker, 2020; Warne et al., 2022; Prescott et al., 2024; Prescott and Baker, 2021). However, these methods inherit limitations of ABC approaches, particularly in high-dimensional parameter spaces, where neural density estimators offer more scalable alternatives to complex real-world problems (Lueckmann et al., 2021). Concurrently with our work, Thiele et al. (2025) developed a multifidelity SBI approach based on response distillation, Hikida et al. (2025) adapted multilevel Monte Carlo techniques to SBI, and Saoulis et al. (2025) applied transfer learning to accelerate inference on a cosmological task.
Beyond SBI, multifidelity has been explored in Bayesian optimization, where Gaussian process models integrate data of different fidelities to infer expensive functions (e.g., Song et al., 2019; Zanjani Foumani et al., 2023). These approaches focus on learning surrogate likelihood functions rather than posteriors over simulator parameters, but they highlight the broad applicability of the multifidelity concept.
@@
To facilitate learning in a target domain, transfer learning borrows knowledge from a source domain (Panigrahi et al., 2021). This is often done when the target dataset is smaller than the source dataset (Larsen-Freeman, 2013). For numerical simulators, transfer learning approaches have been used to lower the simulation budget, for instance, in CO2 forecasting (Falola et al., 2023), surrogate modeling (Wang et al., 2024; Zeng et al., 2026), and model inversion with physics-informed neural networks (Haghighat et al., 2021). To the best of our knowledge, the potential of transfer learning for computationally efficient simulation-based inference has not been fully realized yet.
@@
Recent work reduces the cost of SBI for expensive simulators through active learning or efficient representations. Active learning methods adaptively select simulation parameters for neural likelihood or posterior estimation (Lueckmann et al., 2019; Griesemer et al., 2024), paralleling Bayesian optimization for ABC (Gutmann and Corander, 2016). Efficiency also improves through learned representations such as signature-based features (Dyer et al., 2022), compositional models (Gloeckler et al., 2025), or self-consistency objectives (Schmitt et al., 2024a; b). Unlike these single-fidelity approaches, MF-(TS)NPE leverages an expert-designed low-fidelity simulator and combines transfer learning with active learning to refine posterior estimates efficiently.
@@ 3 Methods
MF-(TS)NPE is a multifidelity approach to Neural Posterior Estimation (NPE) for computationally expensive simulators leveraging transfer learning and, in its sequential variant, active learning. We present our approach in Sec. 3.1. In Sec. 3.1.4, we discuss the evaluation metrics used to compare our method against NPE (Greenberg et al., 2019), TSNPE (Deistler et al., 2022), and MF-ABC (Prescott and Baker, 2020). MF-(TS)NPE is summarized in Fig. 1, Algorithms 1 and 3.
@@ 3.1 Multifidelity NPE
Figure 1: Multifidelity Neural Posterior Estimation proceeds by dense sampling from the prior distribution, running the low-fidelity simulator (e.g., a two-compartment neuron model (Hodgkin and Huxley, 1952)), and training a neural density estimator with a negative log-likelihood loss. MF-NPE then retrains the pre-trained network on sparse samples from the same prior distribution and respective high-fidelity simulations (e.g., a multicompartmental neuron model (Rall, 1995)). Given empirical observations $\boldsymbol{x_{o}}$, MF-NPE estimates the posterior distribution given the high-fidelity model. In the sequential case, the parameters for high-fidelity simulations are drawn from iterative refinements of the prior distribution within the support of the current posterior estimate, at some observation $\boldsymbol{x_{o}}$.
@@
As with NPE (Papamakarios and Murray, 2016; Greenberg et al., 2019), to estimate the posterior density over model parameters $\boldsymbol{\theta}$ for which the likelihood function is unavailable, we consider a sufficiently expressive neural density estimator $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})$, and train it to minimize the negative log-likelihood loss:
@@
where $\boldsymbol{\theta}$ is sampled from the prior distribution, $\boldsymbol{x}$ denotes the respective simulations (i.e., samples from $p(\boldsymbol{x}|\boldsymbol{\theta})$), and $\boldsymbol{\phi}$ are the network parameters. By minimizing $\mathcal{L}(\cdot)$, the neural density estimator approximates the conditional distribution $p(\boldsymbol{\theta}|\boldsymbol{x})$ directly (Papamakarios and Murray, 2016) (proof of convergence in Appendix B). Given an empirical observation $\boldsymbol{x_{o}}$, we can then estimate the posterior over parameters $p(\boldsymbol{\theta}|\boldsymbol{x_{o}})$. To ensure $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x_{o}})$ closely approximates the true posterior $p(\boldsymbol{\theta}|\boldsymbol{x_{o}})$, the density estimator must be sufficiently expressive. We use neural spline flows (NSFs) (Durkan et al., 2019), expressive normalizing flows that have been shown empirically to be competitive for SBI (Lueckmann et al., 2021). To avoid overfitting when training NSFs, we use the same validation-based early stopping criterion $S$ as in the SBI package (Boelts et al., 2024) (details in Appendix.C.1).
@@ 3.1.1 Transfer learning
MF-NPE leverages representations learned from low-fidelity simulations to reduce the number of high-fidelity simulations required to approximate a high-fidelity posterior. To that end, MF-NPE adopts a fine-tuning strategy of transfer learning: Let $\psi$ be the parameters of the low-fidelity neural density estimator $q_{\psi}(\boldsymbol{\theta}|\boldsymbol{x}_{\mathrm{L}})$ and let $\phi$ be the parameters of the high-fidelity density estimator $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})$. MF-NPE minimizes the loss $\mathcal{L(\phi)}=\mathbb{E}_{\theta\sim p(\boldsymbol{\theta})}\mathbb{E}_{x\sim p(\boldsymbol{x}|\boldsymbol{\theta})}\left[-\log q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})\right]$ on the high-fidelity task, where the parameters $\phi$ are initialized on the pretrained low-fidelity network parameters $\psi$. We argue that by pre-training on low-fidelity simulations, the density estimator learns useful features up front (i.e., the feature spaces of the low- and high-fidelity density estimators overlap), so fewer high-fidelity simulations suffice to refine the posterior estimates. Indeed, Tahir et al. (2024) shows that once networks learn suitable features for a given predictive task, they drastically reduce the sample complexity for related tasks. Other strategies to pretraining are discussed in Appendix G.4.
MF-NPE can naturally accommodate more than two fidelity levels (Appendix L), does not require more hyperparameter tuning than NPE (Appendix C.1), and is applicable in situations where the low-fidelity model has fewer parameters than the high-fidelity model. In this setting, the parameters that are exclusive to the high-fidelity model are treated as dummy variables in the pre-trained density estimator. The pre-conditioning with these variables leads to the pre-trained neural density estimator to effectively estimate the prior distribution over the respective parameters (OU3 and OU4 in Appendix I.1). As shown below, our method is compatible with both embedding networks and hand-crafted summary statistics of the observations.
@@
1:Input: NN pairs of $(\boldsymbol{\theta},\boldsymbol{x}_{\text{L}})$; $M$ pairs of ($\boldsymbol{\theta},\boldsymbol{x})$; conditional density estimators $q_{\psi}(\boldsymbol{\theta}|\boldsymbol{x}_{\text{L}})$ and $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})$ with respectively learnable parameters $\psi$ and $\phi$; early stopping criterion $S$.
[…]
3:for epoch in epochs do
4:  train $q_{\psi}$ to minimize $\mathcal{L(\psi)}$ until $S$ is reached.
5:end for
6: Initialize $q_{\phi}$ with weights and biases of trained $q_{\psi}$.  /* High-fidelity model */
[…]
8:for epoch in epochs do
9:  train $q_{\phi}$ to minimize $\mathcal{L(\phi)}$ until $S$ is reached.
10:end for